# Attention as a Hypernetwork

### v3

@@ Abstract
Transformers can under some circumstances generalize to novel problem instances whose constituent parts might have been encountered during training but whose compositions have not. What mechanisms underlie this ability for compositional generalization? By reformulating multi-head attention as a hypernetwork, we reveal that a composable, low-dimensional latent code specifies key-query specific operations. We find empirically that this latent code is predictive of the subtasks the network performs on unseen task compositions revealing that latent codes acquired during training are reused to solve unseen problem instances. To further examine the hypothesis that the intrinsic hypernetwork of multi-head attention supports compositional generalization, we ablate whether making the hypernetwork generated linear value network nonlinear strengthens compositionality. We find that this modification improves compositional generalization on abstract reasoning tasks. In particular, we introduce a symbolic version of the Raven Progressive Matrices human intelligence test which gives us precise control over the problem compositions encountered during training and evaluation. We demonstrate on this task how scaling model size and data enables compositional generalization in transformers and gives rise to a functionally structured latent space.  [^1] Code available at https://github.com/smonsays/hypernetwork-attention
@@ 1 Introduction
Abstract reasoning is a cornerstone of human intelligence. Many intelligence tests in psychology measure intelligence using abstract reasoning tasks [e.g. 1, 2, 3]. Arising from the idea that combinatorial structures are central to human cognitive abilities, the ability of connectionist models to reason has been debated for decades [e.g. 4, 5, 6]. However, with the success of neural networks at scale, the capacity for reasoning has seemingly come into reach [7]. Arguably key to this success is the ability for in-context learning paired with the capacity to flexibly recombine learned skills and knowledge to unseen settings. While these capabilities emerge at least partially from training large models on huge datasets [8], it is not clearly understood how the resulting networks implement them - and why they still often fail.
@@
A number of recent studies that aim to illuminate in-context learning abilities have found that transformers can learn to perform gradient-based optimization within sequence [9, 10, 11]. They build on the insight that linear attention can be understood as a fast-weight programmer [12], that learns to construct an implicit model in-context, with weights given by a sum of input-dependent outer products [13]. Complementing this perspective, [14, 15] have found that in-context learning creates task vectors in the hidden activations of deeper layers of the network which summarize in-context information and modulate the forward computation. What is striking is that learning in-context can under some circumstances lead to compositional generalization, that is generalization to unseen combinations of previously observed constituents [16, 17, 18], a capacity that neural networks notoriously struggle with [19, 20, 21]. The mechanism behind this ability however is not well understood. Here we aim to shed light on this question.
@@
Central to our results is a novel decomposition of multi-head attention as a hypernetwork, a neural network that reconfigures the computation of another neural network in a data-dependent manner. Our decomposition differs from that of [13], revealing that multiple heads form a linear hypernetwork with key-query specific inputs, illustrated in Figure 1. As a result, the attention scores along the head index represent a compact code that specifies the operation applied to each key-query pair via a linear value network. We hypothesize that the shared hypernetwork in each attention layer is what supports the reuse and recombination of learned operations. In order to investigate whether this constitutes a faithful characterization of what transformers learn in practice, we study in-context learning tasks that require recomposing knowledge obtained during training. As one such task we develop a challenging abstract reasoning task based on the Raven Progressive Matrices human intelligence test [3] and show how multi-head attention develops a structured latent code that captures information about the operations implemented.
@@
  • We reformulate standard multi-head attention from a hypernetwork perspective revealing that network computations are specified by a low dimensional latent code.
@@
  • To test the hypothesis that the hypernetwork mechanism supports compositionality, we introduce a simple modification to multi-head linear attention that makes the value network nonlinear without introducing additional parameters finding that it improves compositional generalization.
  • We introduce sraven, a challenging Raven-based symbolic, abstract reasoning task whose difficulty can be controlled parametrically and which offers precise control over the problem compositions encountered during training in order to test for compositional generalization.
@@ 2 Attention as a hypernetwork
In this section, we will first briefly recapitulate hypernetworks [22] and standard multi-head dot-product attention [23] before showing how attention can equivalently be expressed as a hypernetwork. We will then introduce Hypernetwork Linear Attention (HYLA) as a simple modification of linear attention that renders the hypernetwork mechanism inside attention more expressive.
For notation we will use bold lower-case letters for vectors (e.g. $\bm{z}$), bold upper-case letters for matrices (e.g. $\bm{\mathrm{A}}$) and bold, italic upper-case letters for learnable parameter matrices (e.g. $\bm{W}_{V}$).
@@ 2.1 Multi-head attention
A hypernetwork is a neural network $h(\bm{z};\bm{\theta})$ parameterized by $\bm{\theta}$ that takes as input a latent code $\bm{z}$ and outputs parameters $\bm{\mathrm{W}}$ that parameterize a value network $f(\bm{x};\bm{\mathrm{W}})$ which then processes the inputs (see Figure 1A). The latent code, $\bm{z}$, is typically low-dimensional and can be interpreted as a specification of the computation performed by the value network. As we will see in the following, multi-head attention can be viewed as a linear hypernetwork producing the weights for a linear value network. Note that this is different from the observation that linear attention can be understood as a fast weight programmer [13] where the fast weights of a value network are constructed as a sum of outer products over the key indices instead of the head indices.
Self-attention maps a sequence of inputs $\bm{\mathrm{X}}\in\mathbb{R}^{T\times D}$ to a sequence of outputs $\bm{\mathrm{Y}}\in\mathbb{R}^{T\times D}$. For each attention head $h\in\{1,\ldots,H\}$ the inputs are projected into keys $\bm{\mathrm{K}}_{h}=\bm{\mathrm{X}}\bm{W}^{\text{key}}_{h}$ and queries $\bm{\mathrm{Q}}_{h}=\bm{\mathrm{X}}\bm{W}^{\text{query}}_{h}$ using head-specific projection matrices $\bm{W}^{\text{key}}_{h},\bm{W}^{\text{query}}_{h},\bm{W}^{\text{value}}_{h}\in\mathbb{R}^{D\times D^{\text{key}}}$. The attention matrix can then be obtained as
$\bm{\mathcal{A}}=\sigma\bigl{(}\;\left[\bm{\mathrm{\tilde{A}}}_{1}\;\bm{\mathrm{\tilde{A}}}_{2}\;\ldots\;\bm{\mathrm{\tilde{A}}}_{H}\right]\;\bigr{)}\text{ with }\bm{\mathrm{\tilde{A}}}_{h}=\frac{\bm{\mathrm{Q}}_{h}\bm{\mathrm{K}}_{h}^{\top}}{\sqrt{D^{\text{key}}}}$ (1)
where $\bm{\mathcal{A}}=(a_{h,q,k})_{h,q,k}\in\mathbb{R}^{H\times T\times T}$ stacks the head-specific unnormalized attention matrices $(\bm{\mathrm{\tilde{A}}}_{h})_{h}$ into a three-dimensional array and applies a normalization operation $\sigma(\cdot)$. For linear attention, $\sigma(\cdot)=\mathrm{Id}(\cdot)$ is simply the identity whereas for softmax attention $\sigma(\cdot)=\mathrm{Softmax}(\cdot)$ applies the softmax operation for each head and query index independently across the key indices. Considering a single query index $q\in\{1,\ldots,T\}$ and $\mathbf{x}_{k}$ the $k$-th row of $\bm{\mathrm{X}}$, multi-head attention (MHA) can be equivalently expressed as a hypernetwork with a key-query specific latent code:
@@
where $\oplus$ denotes concatenation along the leading axis and the $\bm{W}^{\text{out}}_{h}\in\mathbb{R}^{D^{\text{value}}\times D}$ are defined as the slices of $\bm{W}^{\text{out}}$ such that $\bm{W}^{\text{out}}=\bigoplus_{h=1}^{H}\bm{W}^{\text{out}}_{h}$. See Figure 1C for a diagram of the computation.
As a result, the key-query specific attention scores over the multiple heads form a latent code where the number of heads $H$ corresponds to the latent dimension of a hypernetwork. Accordingly, the dot products between any key-query pair over all heads can be interpreted as an amortized inference process to configure the computation implemented in the linear value network.
The key significance of our finding is that it suggests that multi-head attention can implement specialized, reusable operations through the latent code of the hypernetwork. Notably, it has been shown that hypernetworks support compositional generalization [24], providing a possible explanation for similar capabilities observed in transformers [17, 16, 18]. Note that inside of multi-head attention the same hypernetwork is used for every key-query index. Specifically, the hypernetwork linearly combines the same matrices $(\bm{W}^{\text{out}}_{h},\bm{W}^{\text{value}}_{h})_{h}$ according to the weights given by $a_{h,q,k}$. This incentivizes reuse of latent codes and their associated operations implemented through the value network.
@@ 2.2 Hypernetwork Linear Attention (HYLA)
If the perspective of multi-head attention as a hypernetwork indeed reflects how it operates in practice, it should be possible to modify multi-head attention in a way that further reinforces the hypernetwork mechanism. For instance both the hypernetwork and the value network in multi-head attention are simple linear networks and could be replaced by nonlinear networks. Furthermore the normalization operation $\sigma(\cdot)$ could be used to encode inductive biases such as competition or sparsity in the latent code by imposing structural constraints.
Here, we focus on a simple modification to linear attention where we make the value network nonlinear and normalize the latent code along the head index (instead of along the key index as in standard softmax attention). Specifically, we define Hypernetwork Linear Attention (HYLA) to use a single hidden layer value network with nonlinearity. This should increase the expressivity of the subfunctions that are learned and recomposed by the layer. Noting that the output and value projections, $\bm{W}^{\text{out}}_{h},\bm{W}^{\text{value}}_{h}$, already form a deep linear network in standard multi-head attention (c.f. Eq. 4) this can be achieved without adding additional parameters as follows:
@@
where $\phi(x)$ is an element-wise nonlinearity; here we set it to be a ReLU, $\phi(x)=\max(0,x)$.
Additionally, we use $\sigma(\cdot)=\text{RMSHead}(\cdot)$, to normalize the attention scores for each key and query index independently across the head indices using the $\text{RMSNorm}(\bm{x})=\frac{\bm{x}}{\sqrt{\frac{1}{n}\sum_{i=1}^{n}x_{i}^{2}}}$. This ensures that the parameters of the value network generated by the hypernetwork are appropriately initialized to prevent exploding/vanishing gradients when backpropagating through it. The resulting HYLA-layer is a simple drop-in replacement to standard attention. Notably, since the normalization operation $\sigma(\cdot)$ is local to the query and key index, it requires no communication across keys as opposed to softmax attention. By making the value network more expressive, we allow the network to implement more complex operations for a given latent code which we hypothesize strengthens the ability for compositional generalization.
@@ 3 Compositional generalization on fuzzy logic functions
Figure 2: Compositional generalization on fuzzy logic functions. A We split fuzzy logic functions according to their constituent terms into train and out-of-distribution (OOD) set to measure compositional generalization in a sequence model that learns these functions in-context. B The latent code of the query token is predictive of the constituent terms underlying each task. Shown is the F1 score on unseen tasks of logistic regression classifiers for each layer and term, trained to predict the terms underlying each task based on the attention scores across the head index for the query attending to itself. C Task performance on unseen tasks reported as OOD $R^{2}$ for varying number of in-context examples and fraction of tasks held out during training. D tSNE visualization of the latent codes (attention scores across the head index for the query token attending to itself) colored according to the target label (top) and colored according to the first term of the fuzzy logic function of each task (bottom).
We start by developing a fuzzy logic task with compositional structure, inspired by a similar task previously considered by Rahaman et al. [25]. Learning such tasks in an in-context learning setting will allow us to study the ability of multi-head attention-based models to compositionally generalize and analyze the structure of their latent code when solving each task.
Given a set of $L$ scalars $\{x_{1},...,x_{L}\}$ with $x_{i}\in[0,1]$ we define fuzzy logic operators on the elements of this set as the Zadeh operators [26]:
@@
where each term is a conjunction of all variables with negations applied to some of them. In order to test for compositional generalization, we index each of the $2^{L}$ possible terms, generate all possible combinations of $K$ terms and hold out a fraction of the combinations for out-of-distribution (OOD) testing, ensuring that their constituent terms have been encountered in some tasks during training.
For each task, we present $N$ examples in-context (see Figure 2A) where each token has dimension $L+1$ consisting of the concatenated inputs sampled uniformly from $[0,1]$ and the corresponding target of the function for the current task. For the final token we mask out the target and use the output logits of the final token to predict the target, measuring the loss using the mean-squared error.

### v4

@@ Abstract
Transformers can under some circumstances generalize to novel problem instances whose constituent parts might have been encountered during training, but whose compositions have not. What mechanisms underlie this ability for compositional generalization? By reformulating multi-head attention as a hypernetwork, we reveal that a composable, low-dimensional latent code specifies key-query specific operations. We find empirically that this latent code is predictive of the subtasks the network performs on unseen task compositions, revealing that latent codes acquired during training are reused to solve unseen problem instances. To further examine the hypothesis that the intrinsic hypernetwork of multi-head attention supports compositional generalization, we ablate whether making the hypernetwork-generated linear value network nonlinear strengthens compositionality. We find that this modification improves compositional generalization on abstract reasoning tasks. In particular, we introduce a symbolic version of the Raven’s Progressive Matrices human intelligence test, which gives us precise control over the problem compositions encountered during training and evaluation. We demonstrate on this task how scaling model size and data enables compositional generalization in transformers and gives rise to a functionally structured latent space.  [^1] Code available at https://github.com/smonsays/hypernetwork-attention
@@ 1 Introduction
Abstract reasoning is a cornerstone of human intelligence. Many intelligence tests in psychology measure intelligence using abstract reasoning tasks (e.g. Wechsler, 1958; Binet & Simon, 1961; Raven, 1962). Arising from the idea that combinatorial structures are central to human cognitive abilities, the ability of connectionist models to reason has been debated for decades (e.g. Rumelhart & McClelland, 1986; Fodor & Pylyshyn, 1988; Smolensky, 1991). However, with the success of neural networks at scale, the capacity for reasoning has seemingly come into reach (Huang & Chang, 2023). Arguably key to this success is the ability for in-context learning paired with the capacity to flexibly recombine learned skills and knowledge to unseen settings. While these capabilities emerge at least partially from training large models on huge datasets (Brown et al., 2020), it is not clearly understood how the resulting networks implement them - and why they still often fail.
@@
A number of recent studies that aim to illuminate in-context learning abilities have found that transformers can learn to perform gradient-based optimization within sequence (Dai et al., 2023; Akyürek et al., 2023; von Oswald et al., 2023). They build on the insight that linear attention can be understood as a fast-weight programmer (Schmidhuber, 1992), that learns to construct an implicit model in-context, with weights given by a sum of input-dependent outer products (Schlag et al., 2021). Complementing this perspective, (Hendel et al., 2023; Todd et al., 2024) have found that in-context learning creates task vectors in the hidden activations of deeper layers of the network, which summarize in-context information and modulate the forward computation. What is striking is that learning in-context can under some circumstances lead to compositional generalization, that is generalization to unseen combinations of previously observed constituents (An et al., 2023; Lake & Baroni, 2023; Hosseini et al., 2022), a capacity that neural networks notoriously struggle with (Srivastava et al., 2023; Press et al., 2023; Dziri et al., 2023). The mechanism behind this ability however is not well understood. Here we aim to shed light on this question.
Our key finding is that through the use of multiple heads, the attention mechanism of transformers is mathematically equivalent to a hypernetwork, a neural network that reconfigures the computation of another neural network as shown in Figure 1. This results in a novel interpretation of the attention scores along the head index as a compact latent code that specifies the operations the attention layer applies to each input. Notably, we hypothesize that because multi-head attention shares the same hypernetwork per layer, it is encouraged to reuse and recombine previously acquired operations.
@@
In order to investigate whether this constitutes a faithful characterization of what transformers learn in practice, we study in-context learning tasks that require recomposing knowledge obtained during training. As one such task, we develop a challenging abstract reasoning task based on the Raven’s Progressive Matrices human intelligence test (Raven, 1962) and show how multi-head attention develops a structured latent code that captures information about the operations implemented.
@@
  • We reformulate standard multi-head attention from a hypernetwork perspective, revealing that network computations are specified by a low dimensional latent code.
@@
  • To test the hypothesis that the hypernetwork mechanism supports compositionality, we introduce a simple modification to multi-head linear attention that makes the value network nonlinear without introducing additional parameters, finding that it improves compositional generalization.
  • We introduce sraven, a challenging symbolic, abstract reasoning task whose difficulty can be controlled parametrically and which offers precise control over the problem compositions encountered during training in order to test for compositional generalization.
@@ 2 Attention as a hypernetwork
In this section, we will first briefly recapitulate hypernetworks (Ha et al., 2017) and standard multi-head dot-product attention (Vaswani et al., 2017) before showing how attention can equivalently be expressed as a hypernetwork. We will then introduce Hypernetwork Linear Attention (HYLA) as a simple modification of linear attention that renders the hypernetwork mechanism inside attention more expressive.
For notation, we will use bold lower-case letters for vectors (e.g. $\bm{z}$), bold upper-case letters for matrices (e.g. $\bm{\mathrm{A}}$) and bold, italic upper-case letters for learnable parameter matrices (e.g. $\bm{W}_{V}$).
@@ 2.1 Multi-head attention
A hypernetwork is a neural network $h(\bm{z};\bm{\theta})$ parameterized by $\bm{\theta}$ that takes as input a latent code $\bm{z}$ and outputs parameters $\bm{\mathrm{W}}$ that parameterize a value network $f(\bm{x};\bm{\mathrm{W}})$, which then processes the inputs (see Figure 1A). The latent code $\bm{z}$ is typically low-dimensional and can be interpreted as a specification of the computation performed by the value network. As we will see in the following, multi-head attention can be viewed as a linear hypernetwork producing the weights for a linear value network. Note that this is different from the observation that linear attention can be understood as a fast weight programmer (Schlag et al., 2021), where the fast weights of a value network are constructed as a sum of outer products over the key indices instead of the head indices.
Self-attention [^2] We focus on self-attention, but the same derivation holds for cross-attention, where the keys and values are obtained as the output of an encoder and the queries come from a decoder. maps a sequence of inputs $\bm{\mathrm{X}}\in\mathbb{R}^{D\times T}$ to a sequence of outputs $\bm{\mathrm{Y}}\in\mathbb{R}^{D\times T}$. For each attention head $h\in\{1,\ldots,H\}$, the inputs are projected into keys $\bm{\mathrm{K}}_{h}=\bm{W}^{\text{key}}_{h}\bm{\mathrm{X}}$ and queries $\bm{\mathrm{Q}}_{h}=\bm{W}^{\text{query}}_{h}\bm{\mathrm{X}}$ using head-specific projection matrices $\bm{W}^{\text{key}}_{h},\bm{W}^{\text{query}}_{h},\bm{W}^{\text{value}}_{h}\in\mathbb{R}^{D^{\text{head}}\times D}$ with $D^{\text{head}}=\frac{D}{H}$. The attention matrix can then be obtained as
$\bm{\mathcal{A}}=\sigma\bigl{(}\;\left[\bm{\mathrm{\tilde{A}}}_{1}\;\bm{\mathrm{\tilde{A}}}_{2}\;\ldots\;\bm{\mathrm{\tilde{A}}}_{H}\right]\;\bigr{)}\text{ with }\bm{\mathrm{\tilde{A}}}_{h}=\frac{\bm{\mathrm{Q}}_{h}^{\top}\bm{\mathrm{K}}_{h}}{\sqrt{D^{\text{head}}}}$ (1)
where $\bm{\mathcal{A}}=(a_{h,q,k})_{h,q,k}\in\mathbb{R}^{H\times T\times T}$ stacks the head-specific unnormalized attention matrices $(\bm{\mathrm{\tilde{A}}}_{h})_{h}$ into a three-dimensional array and applies a normalization operation $\sigma(\cdot)$. For linear attention, $\sigma(\cdot)=\mathrm{Id}(\cdot)$ is simply the identity whereas for softmax attention $\sigma(\cdot)=\mathrm{Softmax}(\cdot)$ applies the softmax operation for each head and query index independently across the key indices. Considering a single query index $q\in\{1,\ldots,T\}$ and $\mathbf{x}_{k}\in\mathbb{R}^{D}$ the $k$-th column of $\bm{\mathrm{X}}$, multi-head attention (MHA) can be equivalently expressed as a hypernetwork:
@@
where $\oplus$ denotes concatenation and the $\bm{W}^{\text{out}}_{h}\in\mathbb{R}^{D\times D^{\text{head}}}$ are defined as the slices of $\bm{W}^{\text{out}}\in\mathbb{R}^{D\times D}$ such that $\bm{W}^{\text{out}}=\bigoplus_{h=1}^{H}\bm{W}^{\text{out}}_{h}$. See Figure 1C for a diagram of the computation.
As a result, the key-query specific attention scores over the multiple heads form a latent code, where the number of heads $H$ corresponds to the latent dimension of a hypernetwork. Accordingly, the dot products between any key-query pair over all heads can be interpreted as an amortized inference process to configure the computation implemented in the linear value network.
The key significance of our finding is that it suggests that the use of multiple heads in attention allows it to implement specialized, reusable operations through the latent code of a hypernetwork. It has been theoretically and empirically shown that hypernetworks support compositional generalization (Schug et al., 2024), providing a possible explanation for similar capabilities observed in transformers (Lake & Baroni, 2023; An et al., 2023; Hosseini et al., 2022). Note that multi-head attention reuses the same hypernetwork for every key-query index. Specifically, the hypernetwork linearly combines the same matrices $(\bm{W}^{\text{out}}_{h},\bm{W}^{\text{value}}_{h})_{h}$ according to the weights given by $a_{h,q,k}$. This incentivizes reuse of latent codes and their associated operations implemented through the value network.
@@ 2.2 Hypernetwork Linear Attention (HYLA)
If the perspective of multi-head attention as a hypernetwork indeed reflects how it operates in practice, it should be possible to modify multi-head attention in a way that further reinforces the hypernetwork mechanism. For instance, both the hypernetwork and the value network in multi-head attention are simple linear networks and could be replaced by nonlinear networks. Furthermore, the normalization operation $\sigma(\cdot)$ could be used to encode inductive biases such as competition or sparsity in the latent code by imposing structural constraints.
Here, we focus on a simple modification to linear attention, where we make the value network nonlinear and normalize the latent code along the head index (instead of along the key index as in standard softmax attention). Specifically, we define Hypernetwork Linear Attention (HYLA) to use a single hidden layer value network by adding a nonlinearity. This should increase the expressivity of the subfunctions that are learned and recomposed by the layer. Noting that the output and value projections $\bm{W}^{\text{out}}_{h},\bm{W}^{\text{value}}_{h}$ already form a deep linear network in standard multi-head attention (c.f. Eq. 4) this can be achieved without adding additional parameters as follows:
@@
where $\phi(x)$ is an element-wise nonlinearity; here we set it to be a ReLU, $\phi(x)=\max(0,x)$. Additionally, we use $\sigma(\cdot)=\text{RMSHead}(\cdot)$, to normalize the attention scores for each key and query index independently across the head indices using $\text{RMSNorm}(\bm{x})=\frac{\bm{x}}{\sqrt{\frac{1}{n}\sum_{i=1}^{n}x_{i}^{2}}}$ (without adding a learnable scalar as in Zhang & Sennrich (2019)). This ensures that the parameters of the value network generated by the hypernetwork maintain the variance preserving properties of neural network initializations that have been found to be important for the stability of gradient-based training (Glorot & Bengio, 2010). The resulting HYLA-layer is a simple drop-in replacement to standard attention. Notably, since the normalization operation $\sigma(\cdot)$ is local to the query and key index, it requires no communication across keys as opposed to softmax attention. By making the value network more expressive, we allow the network to implement more complex operations for a given latent code which we hypothesize strengthens the ability for compositional generalization.
@@ 3 Compositional generalization on fuzzy logic functions
Figure 2: Compositional generalization on fuzzy logic functions. A We split fuzzy logic functions according to their constituent terms into train and out-of-distribution (OOD) sets to measure compositional generalization in a sequence model that learns these functions in-context. B The latent code of the response token is predictive of the constituent terms underlying each task. Shown is the F1 score on unseen tasks of logistic regression classifiers for each layer and term, trained to predict the terms underlying each task based on the attention scores across the head index for the response token attending to itself. C Task performance on unseen tasks reported as OOD $R^{2}$ for varying number of in-context examples and fraction of tasks held-out during training. D tSNE visualization of the latent codes (attention scores across the head index for the response token attending to itself) colored according to the target label (top) and colored according to the first term of the fuzzy logic function of each task (bottom).
We start by developing a fuzzy logic task with compositional structure, inspired by a similar task previously considered by Rahaman et al. (2021). Learning such tasks in an in-context learning setting will allow us to study the ability of multi-head attention-based models to compositionally generalize and analyze the structure of their latent code when solving each task.
Given a set of $L$ scalars $\{x_{1},...,x_{L}\}$ with $x_{i}\in[0,1]$ we define fuzzy logic operators on the elements of this set as the Zadeh operators (Zadeh, 1965):
@@
where each term is a conjunction of all variables with negations applied to some of them. In order to test for compositional generalization, we index each of the $2^{L}$ possible terms, generate all possible combinations of $K$ terms and hold out a fraction of the combinations for out-of-distribution (OOD) testing, ensuring that their constituent terms have been encountered in some tasks during training. For each task, we present $N$ examples in-context (see Figure 2A) where each token has dimension $L+1$, consisting of the concatenated inputs sampled uniformly from $[0,1]$ and the corresponding target of the function for the current task. For the final token, we mask out the target and use the output logits of the final token to predict the target, measuring the loss using the mean-squared error.